OpenStudy (anonymous):

the fastest a human has ever run is 27 miles per hour. How many miles per minute did the human run?

OpenStudy (anonymous):

plz help anyone

OpenStudy (anonymous):

all you have to do is find out how much a mile is then multiply it by 27

OpenStudy (austinl):

\(\dfrac{27~\text{miles}}{1~\text{hour}}\times\dfrac{1~\text{hour}}{60~\text{minutes}}=~?\)

OpenStudy (austinl):

The hours cancel and you are left with \(\dfrac{27}{60}\frac{mi}{min}\) Which can be simplified, or possibly depending on the course approximated in decimal form.

OpenStudy (anonymous):

so it = to 27 over 60

OpenStudy (austinl):

But that can be simplified.

OpenStudy (anonymous):

ok to wat

OpenStudy (anonymous):

to wat do u simplify it to

OpenStudy (anonymous):

It can be simplified to 9 over 10

OpenStudy (anonymous):

look 27 is begger than 60 so it can not be simplifyed

OpenStudy (anonymous):

\[(3*3*3)/(2*2*3*5)\]